the 10th term of an arithmetic seq. is 25 and the sum of the 1st 20 terms is 530. find the 17th term of the seq.

Found 2 solutions by srinivas.g, stanbon:
Answer by srinivas.g(540)

10 th term =25
In A.p, formula for n th term = a +(n-1)d
where a = initial or starting term
d= common difference
since 10 th term = 25
a+(10-1)d =25
a+9d =25............. ...(1)
sum of n terms = (first term + last term ) * no of terms / 2

here last term = 20 th term = a+(20-1)d
=a+19 d
sum = (a+(a+19d)) *20 /2
530 = ( 2a+19d ) *10
divide with 10 on both sides
530/10 = (2a+19d) * 10/10
53 = 2a+19d *1
so 2a+19d = 53 ...................(2)
so we need to solve equations (1) and (2) to get 'a ' and ' d'
multiply eq (1) with 2
2*(a+9d) = 2(25)
2a+18d =50 ...............(3)
subtract eq (3) from eq(2)
2a + 19d -(2a+18d) = 53-50
d =3
put d= 3 in eq(1) a+9(3) =25
a= -2
so 17 th term = a+(17-1)d
= -2 +16*3
=-2 +48
=46

Answer by stanbon(75887)

the 10th term of an arithmetic seq. is 25 and the sum of the 1st 20 terms is 530. find the 17th term of the seq.
---
Equations:
a(10) = a + 9d = 25
a(20) = a + 19d
----
S(20) = (20/2)(a + (a+19d)) = 530
10(2a+19d) = 530
-----------------
Equations:
a + 9d = 25
2a+19d = 53
-----
Modify:
2a + 18d = 50
2a + 19d = 53
---
Subtract to get:
d = 3 (the common difference)
Solve for "a":
a + 9d = 25
a + 27 = 25
a = -2
===================
Seventeenth term:
a(17) = a + 16d
a(17) = -2 + 16*3 = 46
-------------------------
Cheers,
Stan H.
====================